What volume of 0.125 M CaCI_2 contains 2.22 times
Answer #1

Cl- moles = number of Cl- ions / Avagadro number

            = ( 2.22 x 10^21) / ( 6.023 x 10^23 ) = 0.003686

CaCl2 moles = ( 1/2) CaCl2 moles = ( 1/2) (0.003686) = 0.001843   ( since 1CaCl2 gives 2Cl-)

Molarity of CaCl2 = Moles of CaCl2 / Volume of CaCl2

0.125 M = 0.001843 / vol

vol = 0.01474 L = 14.74 ml